At a point 3200 km vertically above the surface of the earth, acceleration due to gravity
of earth in SI units is
a) 6.66 b) 3.33 c) 5.55 d) 4.44

Grade:

2 Answers

Souvik Banerjee
7 years ago
radius of earth = 6.38 *10^6 m. value of G=6.67*10^-11 Nm^2Kg^-2.
distance of the point =3200km=3.2*10^6m. mass of earth M =5.97*10^24 kg.
new radius (r) = 6.38*10^6+3.2*10^6.
g=GM/r^2 (on substituting the values of G M and r ) we get g=4.34m/s^2
Ragul Krishna
one year ago
At a point 3200 km vertically above the surface of the earth, acceleration due to gravity
of earth in SI units